MICROWAVE SPECTRA OF O-BENZYNE PRODUCED IN A DISCHARGE NOZZLE ^{a}

\underline{STEPHEN G. KUKOLICH}, CHAKREE TANJAROON, Department of Chemistry, University of Arizona, Tucson, AZ 85721 ; MICHAEL C. McCARTHY, and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics and Division of Applied Sciences, Harvard University, Cambridge, MA 02138.

Benzynes have been known as reactive intermediates in organic reactions for many years, and have recently been implicated in gasoline combustion reactions and antitumor activity of enediyenes. Rotational transitions were measured for the normal isotopomer, the two unique single-D isotopomers and the ^{13}C_{6} isotopomer of o-benzyne(o-didehydrobenzene), using a pulsed-beam Fourier transform spectrometer with an electric discharge nozzle source, using a dilute(<1%) mixture of benzene in neon. 27 b-dipole transitions for the normal isotopomer were measured in the 7-40 GHz range to obtain A = 6989.7303(72), B = 5706.8045(71) and C = 3140.3714(50) MHz, and five centrifugal distortion constants. The inertial defect is Δ = 0.06935 amu Å^{2}, consistent with a planar structure. Deuterium hyperfine structure for the D_{1}(closest to the C-C triple bond) and D_{2} (furthest from the triple bond), was analysed to obtain deuterium quadrupole coupling strengths eQq_{aa} = 231(17) kHz, eQq_{bb} = -88(13) kHz, for D_{1}, and eQq_{aa} = 29(12), and eQq_{bb} = 83(12) kHz, for D_{2}. The C-D bond axis quadrupole coupling strengths will be compared with values for benzene.

aSupported by THE NATIONAL SCIENCE FOUNDATION